OpenStudy (calculusxy):

Jack uses 10 gallons of water on his garden on a partly cloudy day. This is 3/4 of the amount of water he uses on his garden on a sunny day. How many gallons of water does Jack use on his garden on a sunny day?

OpenStudy (calculusxy):

@wio

OpenStudy (anonymous):

The equation here is: \[ x \times \frac 34 = 10 \]and you need to solve for \(x\).

OpenStudy (calculusxy):

Is the answer 7 1/2 gallons?

OpenStudy (anonymous):

No. Remember he should water the garden MORE on a sunny day than on a cloudy day, and he already does 10 gallons on a cloudy day.

OpenStudy (calculusxy):

okay to solve for x i should do: \[\frac{ 10 }{ 1 } \div \frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

SOrry i meant to mutiply

OpenStudy (calculusxy):

is it 13 1/3 gallons?

OpenStudy (calculusxy):

that should equal to 10 right when i do 40/3 * 3/4?

OpenStudy (calculusxy):

\[\frac{ 120 }{ 12 }\]

OpenStudy (anonymous):

yeah, \(120/12 = 10\)

OpenStudy (calculusxy):

help on the other question plz